Problems 5 and 6: Endpoint A and midpoint M of are given. Find endpoint B.
5. A(1,2) and M(5,8). Answer: B(9,14). Show all the solution steps in detail, using the example as a model.
6. A(-2,-1) and M(0,0).

how do I figure it out?

Answer by stanbon(75887)

Endpoint A and midpoint M of are given. Find endpoint B.
5. A(1,2) and M(5,8). Answer: B(9,14). Show all the solution steps in detail, using the example as a model.
-----
Let the end point be (x,y)
----
Equation:
(1+x)/2 = 5
1+x = 10
x = 9
----
(2+y)/2 = 8
2+y = 16
y = 14
------------------
6. A(-2,-1) and M(0,0).
Let the end point be (x,y)
Equations:
(x-2)/2 = 0
x-2 = 0
x = 2
----
(y-1)/2 = 0
y-1 = 0
y = 1
=============
Cheers,
Stan H.
=============